OpenStudy (anonymous):

For the equation x2 + 3x + j = 0, find all the values of j such that the equation has two real number solutions. Show your work.

OpenStudy (anonymous):

you need \(b^2-4ac>0\) so in this case you must solve \(9-4\times j>0\)

OpenStudy (cwrw238):

for 2 real roots 3^2 - 4*j > 0

OpenStudy (hoblos):

Δ = 3^2 - 4j = 9-4j the equation has two real number solutions for Δ>0 so 9-4j>0 j< 9/4